OpenStudy (anonymous):

If f(x) = 8 - 3|x + 2| find f(-6).

OpenStudy (anonymous):

f(-6) just means that x = -6. can you solve it now?

OpenStudy (anonymous):

so it will be 4

OpenStudy (anonymous):

almost. remember that those vertical lines around the x+2 means to take the absolute value, in other words the positive value of what is inside those lines

OpenStudy (anonymous):

so it will be positive 4 in stead of -4 because of the absolute value

OpenStudy (anonymous):

so 8 - 3|x + 2| where x = -6 8 - 3|-6+2| 8 - 3|-4| the absolute value of -4 = 4, and remember the order of operations, you multiply first. 8 - 3*4 8 - 12 = -4

OpenStudy (anonymous):

ok i got it now